# IDEA: Enhancing the Rule Learning Ability of Large Language Model Agent through Induction, Deduction, and Abduction

### v4

@@ A.1 Figures
Figure 7: Although agents continuously refine their hypotheses toward the ground truth rule, identifying the exact rule remains a challenging task. According to our evaluation, humans have a 43.3% success rate in finding the ground truth rule, with 13% of these discoveries occurring during the abduction stage and 30% during the induction stage. In contrast, LLM agents exhibit a different pattern. They successfully identify the ground truth rule in approximately 30% of puzzles, with nearly all of these discoveries occurring during the abduction stage and only 3% achieved through interaction and induction. This highlights a significant limitation of current LLM agents, as they lack the ability to learn effectively from interactions. Consequently, the rule-learning patterns of LLM agents differ markedly from those of humans.
@@ A.2 Tables
Puzzle ID No. of Functions No. of Unknown Paramters No. of Elementary Functions
1–5 1 1 1
6–10 2 2 2
11–15 2 2 3
16–20 3 3 4
Table 5: Escape room puzzle distribution, the number of paintings ranges from 3 to 13, increasing by one for each subsequent puzzle ID within each group. The first group (1–10) presents all paintings initially, whereas the second group (11–20) requires agent actions to reveal all paintings.
Puzzle ID Number of Paintings Paintings Visibility
1–10 3–13 All paintings visible at start
11–20 3–13 Requires actions to view all paintings
Table 6: Reactor puzzle distribution, five unique target strings need to be synthesized (ACB, CCADD, CADEA, ABCDEF, FEADE) are uniformly distributed across each rule category.
Puzzle ID Reactor Rule
1–5 Simple Concatenation
6–10 Reverse Concatenation
11–15 Middle Insertion
16–20 Prefix Replacement
@@
Deduction Only Gemma-7B 8.28 3.75 11.49 9.59
Llama-8B 2.39 3.4 1.2 2.56
Llama-70B 2.38 2.91 0.44 3.8
GPT-3.5-Turbo 6.18 8.79 5.26 4.5
GPT-4o 1.75 1.55 0.34 3.35
Baseline Gemma-7B 7.27 0.61 11.73 9.46
Llama-8B 3.05 3.38 1.99 3.77
Llama-70B 2.47 1.81 0.92 4.68
GPT-3.5-Turbo 5.97 7.16 2.64 8.12
GPT-4o 2.03 1.5 0.11 4.48
IDEA Gemma-7B 2.74 6.33 1.88 0.02
Llama-8B 1.83 2.29 1.64 1.56
Llama-70B 1.01 0.45 0.39 2.18
GPT-3.5-Turbo 4.49 4.14 2.4 6.93
GPT-4o 1.12 0.52 0.15 2.68
@@ A.3.3 Pseudocode of interactive rule learning procedure
7:     $\mathbb{\tilde{M}}.\text{add}(\text{"You established a new {H} and {P}."})$” )
@@
27:              $\mathbb{\tilde{M}}.\text{add}(\text{"You established a new {H} and {P}."})$” )
@@
A.4 Prompt example
A.4.1 Function Operator Puzzles
@@
A.4.2 Escape Room Puzzles
@@ A.4.3 Reactor Puzzles
A.4.4 Hullicination examples

### v5

@@ A.1 Figures
Figure 7: Although agents continuously refine their hypotheses toward the ground truth rule, identifying the exact rule remains a challenging task. According to our evaluation, humans have a 43.3% success rate in finding the ground truth rule, with 13% of these discoveries occurring during the abduction stage and 30% during the induction stage. In contrast, LLM agents exhibit a different pattern. They successfully identify the ground truth rule in approximately 30% of puzzles, with nearly all of these discoveries occurring during the abduction stage and only 5% achieved through interaction and induction. This highlights a significant limitation of current LLM agents, as they lack the ability to learn effectively from interactions. Consequently, the rule-learning patterns of LLM agents differ markedly from those of humans.
@@ A.2 Tables
No. of Functions No. of Paramters No. Terms No. of Elementary Functions No. of puzzles
1 1 1 1 25
2 2 2 2 25
2 2 3 2 25
3 3 5 4 15
3 4 6 5 10
Table 5: Distribution of puzzles in the Escape Room scenario. For each number of paintings from 3 to 13, there are two visibility conditions: one where all paintings are initially visible, and one where the agent must take actions to reveal all paintings. Under each condition, there are five unique puzzles, resulting in a total of 10 puzzles per number of paintings.
No. of Paintings Visibility Condition No. of Puzzles
3 All visible at start 5
3 Requires actions to reveal 5
4 All visible at start 5
4 Requires actions to reveal 5
13 All visible at start 5
13 Requires actions to reveal 5
Table 6: Distribution of Reactor puzzles across the four rule categories. Each category contains 25 puzzles, drawn from the same set of 25 distinct letter strings, which vary in length from 3 to 6 characters. Each puzzle requires the agent to synthesize a target string according to the specified rule.
Reactor Rule No. Initial Letters No. of Puzzles
Simple Concatenation 2, 3, 4, 5, 6 6, 8, 7, 2, 2
Reverse Concatenation 2, 3, 4, 5, 6 6, 8, 7, 2, 2
Middle Insertion 2, 3, 4, 5, 6 6, 8, 7, 2, 2
Prefix Replacement 2, 3, 4, 5, 6 6, 8, 7, 2, 2
@@ 13 Requires actions to reveal 5
Deduction Only Gemma-7B 6.54 4.05 8.74 6.83
Llama-8B 4.91 3.39 2.85 8.49
Llama-70B 3.39 2.47 1.44 6.25
GPT-3.5-Turbo 8.06 7.52 6.27 10.38
GPT-4o 2.51 2.01 0.65 4.86
Baseline Gemma-7B 7.39 6.12 8.05 8.01
Llama-8B 6.26 6.41 3.24 9.13
Llama-70B 3.36 1.25 1.59 7.23
GPT-3.5-Turbo 6.87 6.85 4.09 9.66
GPT-4o 2.68 1.86 0.19 5.99
IDEA Gemma-7B 5.0 3.65 7.1 4.26
Llama-8B 3.77 3.92 2.73 4.65
Llama-70B 1.73 0.43 0.72 4.05
GPT-3.5-Turbo 5.67 4.69 3.55 8.76
GPT-4o 2.37 1.32 1.25 4.53
@@ A.3.3 Pseudocode of interactive rule learning procedure
7:     $\mathbb{\tilde{M}}.\text{add}(\text{"You established a new {H} and {P}."})$" )
@@
27:              $\mathbb{\tilde{M}}.\text{add}(\text{"You established a new {H} and {P}."})$" )
@@
A.4 Computational Budget
For each setting (the Oracle agent, the Baseline, and the IDEA agent), we ran 300 different puzzles, with costs totaling \$188.10 for GPT-4o and \$37.62 for GPT-3.5. For open-source models, we utilized 8 RTX-A6000 GPUs. LLama3-70B required 5 days to complete its tasks, while Llama3-8b and Gemma-7b each took 2.5 days. We recruited 50 students as human participants, each working for one hour and receiving \$15 as compensation. In total, we spent over \$1,000 and utilized 80 GPU days for open-source models. The experimental cost is notably higher compared to traditional QA datasets due to each puzzle requiring over 20 steps and the rapid accumulation of context during solving.
A.5 Prompt example
A.5.1 Function Operator Puzzles
@@
A.5.2 Escape Room Puzzles
@@ A.5.3 Reactor Puzzles
A.6 Human participants interface
Figure 18: An example of what human participants see during the experiment. Human participants view the same prompt as the LLM agents, except for additional overview information clarifying their responsibilities in the project. The name "{Nick_name}" is a pseudonym used to differentiate participants; we require participants to use only fictitious nicknames rather than their real names to ensure that the data collected is de-identified.
A.6.1 Hullicination examples